OpenStudy (anonymous):

9/(2x+6)-7/(5x+15)=2/3

OpenStudy (mimi_x3):

\[\frac{9}{2x+6}-\frac{7}{5x+15} =\frac{2}{3} \] \[\frac{9*5}{2(x+3)*5}-\frac{7*2}{5(x+3)*2} = \frac{2}{3} \] \[\frac{45}{10(x+3)} -\frac{14}{10(x+3)} =\frac{2}{3} \] \[\frac{31}{10(x+3)} =\frac{2}{3} \] \[2(10x+30) = 31*3\] \[20x+60 = 93\] \[20x=33\] \[x=..\]